An electron has an uncertainty in its position of 609 pm

What is the uncertainty in its velocity?

***Express your answer with the appropriate units.**
Answer #1

use Heisenberg uncertainty principle,

Δx*Δv = h/(4*pi*m)

Here:

Δv = 709 pm = 609*10^-12 m = 6.09*10^-10 m

mass, m = 9.11*10^-31 Kg

Putting values,

6.09*10^-10*Δ v = (6.626*10^-34)/(4*3.14*9.11E-31)

6.09*10^-10*Δ v = 5.791*10^-5

Δv = 9.509*10^4 m/s

Answer: 9.51*10^4 m/s